# Good Things Come in Pairs: Paired Autoencoders for Inverse Problems

### v1

@@ 4.2 Image Restoration: Variational Latent Mapping
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 9: Visualization of paired autoencoders with variational mapping between latent spaces. The leftmost column visualizes the original image $x$ followed by the corrupted image $y$. The next panel shows the mean of the latent distribution computed by the variational encoder, then passed through the decoder. Finally, the rightmost column shows the pixel-wise variance of the reconstructions, computed across $100$ decoded samples from $M^{\dagger}(z_{y})$.
@@ 4.3 Seismic data inversion
In this experiment, the observed seismic wavefield data consist of $63$ receivers, $30$ sources, and $512$ time samples for each of the 48,000///6,000 training/validation examples of size $700\rm{m}\times 700\rm{m}$ from the OpenFWI CurveFaultA dataset [31]. All sources and receivers are located near the top of the computational domain. We used DeepWave [103] to generate seismic data.
@@
Therefore, researchers almost always add at least some random (Gaussian) noise to the synthetic data before solving the inverse problem. However, the magnitude of the noise varies wildly in deep-learning studies for seismic FWI. The authors in [128] use a Signal-to-Noise-Ratio (SNR) of $12$dB and 00dB for different experiments, while [132] report SNRs of 00dB and $6$dB. [52] use a (Peak-Signal-to-Noise-Ratio) PSNR ranging approximately between $39$ and $84$dB. Yet, other studies do not report any noise [26, 125]. Here, we take a somewhat middle ground and add zero-mean Gaussian noise (SNR of $30$dB / PSNR $\approx 35$dB) to the synthetic data for training and we validate using various noise levels between 00dB and $70$dB. Figure 11 shows some samples of the data and its frequency spectra to illustrate that the noise strongly affects the lowest $\lessapprox 3$Hz and highest frequencies.
Figure 11: Three time series and their corresponding amplitude spectra for the noise level in the training set(SNR=$30$dB). Validation experiments use various noise levels, see Figures 13 and 14.
The paired autoencoders provides an approach similar to Equation 8 by using a model decoder $d_{x}(z_{x},\theta_{x}^{\rm{d}})$ to parameterize the inverse problem. What sets the paired autoencoder apart from obtaining a parameterization from a regular VAE decoder or other generative models, is the availability of an accurate initial guess in the latent space, (29), which we use to both warm start and regularize the refinement problem from Equation 28. See Appendix A for the network designs used. Numerical experiments show that the decoded refinement result (30) reduces the data misfit $\|F(\hat{x})-y\|$ and the QoI error in terms of structural similarity index (SSIM) and relative error (RE) $||\hat{x}-x_{\text{true}}||_{2}/||x_{\text{true}}||_{2}$. Figure 12 and Table 2 show the merit of latent-space inversion with an accurate initial guess for a single validation model.
@@
To further investigate the effect of noise on the performance of the direct estimate from the paired autoencoders (27), and the subsequent LSI (28), we repeat the above experiment for various noise levels. The autoencoders are the same as above: trained using data with an SNR of $30$. Figure 13 shows that direct prediction is somewhat impacted by higher noise levels, but subsequent LSI shows to be largely immune to the noise levels selected for this example.
@@
Figure 14: A few examples for validation examples with SNR $=10$dB seismic data. The networks were trained using $30$dB data.
@@ 4.4 Variational Paired Autoencoders for Seismic Full-Waveform Inversion
Figure 15 illustrates how the VPAE supports the objective outlined in the previous paragraph and displays eight samples as defined in Equation 45. For sampling and uncertainty quantification to be practically useful, we need at least a qualitative correlation between the uncertainty and an error estimate in the predictions. Figure 16 shows the summary computed from $100$ samples that highlights a strong correlation between the standard deviation and error in the mean estimate. See Figure 17 for some more examples of the mean, standard deviation, and error computed for the validation examples. All experiments in this subsection use networks trained using data with SNR$=30$dB, while the validation data SNR$=10$dB.
@@ Acknowledgment.
Appendix A Appendix: Neural Network Designs
A.1 Image Restoration
Identical encoder/decoder architectures were used for both autoencoders in the image restoration problems using $28\times 28$ input images. We implement a 6-layer 2D-CNN for our encoder with $3\times 3$ kernels while maintaining a padding of $1$. The first layer increases the number of channels to $16$ and subsequent layers double the number of channels, while halving the spatial dimensions. Here, the spatial downsampling occurs through average pooling with a kernel size and stride of 2. The decoder is implemented using a similar architecture, starting with $128$ channels which are mapped to a $128\times 8\times 8$ tensor via a linear mapping. From here, each block doubles the spatial dimension at each decoder layer while halving the number of channels. Since we upsample by 2×2$\times$2 × at each decoder layer, an additional interpolation is required and is implemented using a nearest-neighbors interpolation algorithm. The full decoder block sequence is given as the following composition of layers
[…]
@@ 20 $1\times 28\times 28$ Sigmoid
A.2 Variational Latent Mapping for Image Restoration
The encoder/decoder architecture used for the probabilistic latent mapping experiment are identical to section A.1. The probabilistic encoder and decoder used for $M^{\dagger}$ is summarized in Table 4. Here we note that the encoder is non-compressive–the value sampled from the probabilistic encoder has dimension $128$ which is equivalent to the dimension of the data latent space.
Image Restoration Encoder
@@ A.2 Variational Latent Mapping for Image Restoration
1 $128$ Linear (hidden dimension $=128$)
@@ 3 $2\times 128$ ReLU
Image Restoration Decoder
@@
1 $128$ Linear (hidden dimension $=128$)
@@ 13-15 $256\times 36\times 32$ Equation 51
16-18 $256\times 18\times 16$ Equation 51 ×$\times 3$
@@ 1 $128\times 9\times 18$ $\mathbf{X}_{i+1}=\mathbf{A}\mathbf{X}_{i}\rightarrow$ reshape
2-4 $256\times 18\times 16$ Equation 51 ×$\times 3$
@@ 7-9 $64\times 18\times 18$ Equation 51
10-12 $64\times 18\times 18$ Equation 50 ×$\times 3$
@@ 2 $64\times 18\times 18$ $\mathbf{X}_{i+1}=\mathbf{K}_{cc}\mathbf{X}_{i}$
3-5 $64\times 18\times 18$ Equation 50 ×$\times 3$

### v2

@@ 4.2 Image Restoration: Variational Latent Mapping
Figure 9: Visualization of paired autoencoders with variational mapping between latent spaces. The leftmost column visualizes the original image $x$ followed by the corrupted image $y$. The next panel shows the mean of the latent distribution computed by the variational encoder, then passed through the decoder. Finally, the rightmost column shows the pixel-wise variance of the reconstructions computed across $100$ decoded samples from $M^{\dagger}(z_{y})$ paired with a colorbar indicating the scale of the variance. All other images are normalized.
@@ 4.3 Seismic data inversion
In this experiment, the observed seismic wavefield data consist of $63$ receivers, $30$ sources, and $512$ time samples for each of the 48,000//6,000 training/validation examples of size $700\rm{m}\times 700\rm{m}$ from the OpenFWI CurveFaultA dataset [35]. All sources and receivers are located near the top of the computational domain. We used DeepWave [117] to generate seismic data.
@@
Therefore, researchers almost always add at least some random (Gaussian) noise to the synthetic data before solving the inverse problem. However, the magnitude of the noise varies wildly in deep-learning studies for seismic FWI. The authors in [146] use a Signal-to-Noise-Ratio (SNR) of 1212dB and 0dB for different experiments, while [151] report SNRs of 0dB and 66dB. [59] use a (Peak-Signal-to-Noise-Ratio) PSNR ranging approximately between $39$ and 8484dB. Yet, other studies do not report any noise [29, 143]. Here, we take a somewhat middle ground and add zero-mean Gaussian noise (SNR of 3030dB / PSNR $\approx 35$dB) to the synthetic data for training and we validate using various noise levels between 0dB and 7070dB. Figure 11 shows some samples of the data and its frequency spectra to illustrate that the noise strongly affects the lowest $\lessapprox 3$Hz and highest frequencies.
Figure 11: Three time series and their corresponding amplitude spectra for the noise level in the training set(SNR=3030dB). Validation experiments use various noise levels, see Figures 13 and 14.
[…]
@@
To further investigate the effect of noise on the performance of the direct estimate from the paired autoencoders (31), and the subsequent LSI (32), we repeat the above experiment for various noise levels. The autoencoders are the same as above: trained using data with an SNR of $30$. Figure 13 shows that direct prediction is somewhat impacted by higher noise levels, but subsequent LSI shows to be largely immune to the noise levels selected for this example. Figure 14 shows another two validation examples using data that contain more noise compared to training data.
@@
Figure 14: A few validation examples with SNR =10=10dB seismic data. The networks were trained using 3030dB data.
@@ 4.4 Variational Paired Autoencoders for Seismic Full-Waveform Inversion
Figure 15 illustrates how the VPAE supports the objective outlined in the previous paragraph and displays eight samples as defined in 49. For sampling and uncertainty quantification to be practically useful, we need at least a qualitative correlation between the uncertainty and an error estimate in the predictions. Figure 16 shows the summary computed from $100$ samples that highlights a strong correlation between the standard deviation and error in the mean estimate. See Figure 17 for some more examples of the mean, standard deviation, and error computed for the validation examples. All experiments in this subsection use networks trained using data with SNR=30=30dB, while the validation data SNR=10=10dB.
@@ Acknowledgment.
6 Neural Network Designs
6.1 Image Restoration
Identical encoder/decoder architectures were used for both autoencoders in the image restoration problems using $28\times 28$ input images. We implement a 6-layer 2D-CNN for our encoder with $3\times 3$ kernels while maintaining a padding of $1$. The first layer increases the number of channels to $16$ and subsequent layers double the number of channels, while halving the spatial dimensions. Here, the spatial downsampling occurs through average pooling with a kernel size and stride of 2. The decoder is implemented using a similar architecture, starting with $128$ channels which are mapped to a $128\times 8\times 8$ tensor via a linear mapping. From here, each block doubles the spatial dimension at each decoder layer while halving the number of channels. Since we upsample by $2\times$ at each decoder layer, an additional interpolation is required and is implemented using a nearest-neighbors interpolation algorithm. The full decoder block sequence is given as the following composition of layers
$\mathbf{X}_{i}\rightarrow{\rm Upsampling\;by\;}2\times\rightarrow{\rm Conv2D}\rightarrow{\rm Batch\;Norm}\rightarrow{\rm SiLU}\rightarrow{\rm Conv2D}\rightarrow\mathbf{X}_{i+1}.$ (53)
@@ 20 $1\times 28\times 28$ Sigmoid
6.2 Variational Latent Mapping for Image Restoration
The encoder/decoder architecture used for the probabilistic latent mapping experiment are identical to Appendix 6.1. The probabilistic encoder and decoder used for $M^{\dagger}$ is summarized in Table 4. Here we note that the encoder is non-compressive–the value sampled from the probabilistic encoder has dimension $128$ which is equivalent to the dimension of the data latent space.
Image Restoration Latent Mapping Encoder
@@ 6.2 Variational Latent Mapping for Image Restoration
1 $128$ Linear (hidden dimension =128=128)
@@ 3 $2\times 128$ ReLU
Image Restoration Latent Mapping Decoder
@@
1 $128$ Linear (hidden dimension =128=128)
@@ 13-15 $256\times 36\times 32$ Equation 55
16-18 $256\times 18\times 16$ Equation 55 $\times 3$
@@ 1 $128\times 9\times 18$ $\mathbf{X}_{i+1}=\mathbf{A}\mathbf{X}_{i}\rightarrow$ reshape
2-4 $256\times 18\times 16$ Equation 55 $\times 3$
@@ 7-9 $64\times 18\times 18$ Equation 55
10-12 $64\times 18\times 18$ Equation 54 $\times 3$
@@ 2 $64\times 18\times 18$ $\mathbf{X}_{i+1}=\mathbf{K}_{cc}\mathbf{X}_{i}$
3-5 $64\times 18\times 18$ Equation 54 $\times 3$